relation between molarity&molality?

Grade:10

2 Answers

TANAYRAJ SINGH CHOUHAN
11 years ago
Relation ship between Molarity (M) and Molality (m) let moles os solid = nB M = nB / V (in ml) x 1000 D(soln) = mass (soln) / V (soln) Mass of solution = Density x Volume d x 1000 [ Let assume of soln = 1000ml ] M = nB / V (in ml) x 1000 M = WB / M wtB x 1000 / 1000 [ M = nB / 1000 x 1000 = nB ] M = WB / MwtB WB = M x MwtB wt of solvent = wt of soln -- wt of solute = 1000d -- M . MwtB Molality (m) = nB / WA (gm) x 1000 m = M / 1000d -- M x MwtB x 1000 ( M = No. of moles of solute ) Here, A = Solvent B = Solute Plz. approve my answer by clicking Yes given below, if u are satisfied by it... Plz/.... And best of luck for your preparation!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
Theresh Babu Benguluri
11 years ago
 
                Molarity/Molality = moles of solute/volume of solution × mass of solvent/moles of solute
 
 
                 = mass of solvent (in kg)/volume of solution (in litre) = mass of solvent (in grams)/volume of solution = W/V
 
 
Let the density of the solution be d g mL−1
 
Mass of solution = V × d
 
Mass of solute = number of moles × molecular mass of solute
 
                = n mA
 
Mass of solvent, W = mass of solution – mass of solute
 
                 = V × d – n × mA
 
 
Thus,         Molaity/Molality (M/m) = V×d–n×mA/V
 
 
                   Molaity (m) = molarity×V/V×d–n×mA
 
 
                                       = molaity/d–n/V×mA
 
                                   m = molaity/d–molaity×mA                                        THERESH